# FlexQ: Efficient Post-training INT6 Quantization for LLM Serving via Algorithm-System Co-Design

### v1

@@ 6 Conclusion
This paper introduces FlexQ, an algorithm-system co-designed framework for LLM inference tailored to INT6 quantization. At the algorithmic level, we integrate fine-grained group quantization with quantization-sensitive layer mixed-precision strategies, achieving negligible accuracy loss under 6-bit quantization. At the system level, we develop the FlexQ W6Ax kernel, which exploits BTCs and introduces memory-efficient bit-level data layouts alongside a suite of novel optimization techniques. FlexQ can be seamlessly integrated into existing inference systems, enabling end-to-end 6-bit support for quantized LLM inference and offering a superior trade-off between model quality and inference efficiency. Extensive experiments demonstrate that FlexQ achieves accuracy comparable to FP16 baselines. In addition, the FlexQ kernel delivers up to 1.82× and 1.27× throughput improvements over cuBLAS and ABQ, respectively, while reducing end-to-end inference latency by 1.33× compared to the state-of-the-art baseline, SmoothQuant. These results validate FlexQ as a practical and highly efficient solution for low-bit quantized LLM inference.

### v2

@@ 6 Conclusion
This paper introduces FlexQ, an algorithm-system co-designed framework for LLM inference tailored to INT6 quantization. At the algorithmic level, we integrate fine-grained group quantization with quantization-sensitive layer mixed-precision strategies, achieving negligible accuracy loss under 6-bit quantization. At the system level, we develop the FlexQ W6Ax kernel, which exploits BTCs and introduces memory-efficient bit-level data layouts alongside a suite of novel optimization techniques. FlexQ can be seamlessly integrated into existing inference systems, enabling end-to-end 6-bit support for quantized LLM inference and offering a superior trade-off between model quality and inference efficiency. Extensive experiments demonstrate that FlexQ achieves accuracy comparable to FP16 baselines. In addition, the FlexQ kernel delivers up to 1.82× and 1.27× throughput improvements over cuBLAS and ABQ-LLM, respectively, while reducing end-to-end inference latency by 1.33× compared to the state-of-the-art baseline, SmoothQuant. These results validate FlexQ as a practical and highly efficient solution for low-bit quantized LLM inference.